# Quantitative Contraction Rates for McKean-Vlasov Stochastic Differential Equations with Multiplicative Noise

### v1

@@ 3.2 Contraction Result Under Dissipative Condition
$-I_{\{r_{t}>0\}}\frac{4Mf^{\prime}(r_{t})}{D}\left(|\alpha_{t}^{\top}e_{t}|^{2}\mathrm{rc}^{2}(U_{t})+|\Delta_{t}^{\top}e_{t}|^{2}\mathrm{sc}^{2}(U_{t})\right)\mathrm{d}t$
@@
The last inequality follows because if $\lambda/2\leq 2L\varepsilon$, then $-\frac{\lambda}{2}G(X_{t},Y_{t})+\frac{\lambda}{2}-2L\varepsilon\leq-\frac{\lambda}{2}G(X_{t},Y_{t})$, and if $22L\varepsilon\leq\lambda/$ 2, then
@@
Furthermore, we notice that using the Cauchy-Schwarz inequality, (1.3), and the fact that $\|H\|=1$, we get
@@
$\leq 2\varepsilon f^{\prime}(r_{t})\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|M(|X_{t}|+|Y_{t}|)$
$\leq\frac{Mf^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t}),$
@@
$\leq Mf^{\prime}(r_{t})\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|G(X_{t},Y_{t})$
$\leq M\sqrt{L_{2}}f^{\prime}(r_{t})\mathrm{sc}^{2}(U_{t})r_{t}G(X_{t},Y_{t})$
$\leq M\sqrt{L_{2}}\delta G(X_{t},Y_{t}).$
@@
$\mathrm{d}[f(r),G(X,Y)]_{t}$ $\leq\frac{Mf^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t})\mathrm{d}t+M\sqrt{L_{2}}\delta G(X_{t},Y_{t})\mathrm{d}t.$ (3.47)
@@
$\quad+G(X_{t},Y_{t})\left[\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+M\sqrt{L_{2}}\right)\delta\right]\mathrm{d}t$
$\quad+G(X_{t},Y_{t})\left[-\frac{4Mf^{\prime}(r_{t})}{D}\left(|\alpha_{t}^{\top}e_{t}|^{2}\mathrm{rc}^{2}(U_{t})+|\Delta_{t}^{\top}e_{t}|^{2}\mathrm{sc}^{2}(U_{t})\right)\right]\mathrm{d}t$
@@
$\quad+\frac{Mf^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t})\mathrm{d}t+\mathrm{d}M_{t}$
@@
$\quad+G(X_{t},Y_{t})\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+M\sqrt{L_{2}}\right)\delta\mathrm{d}t$
@@
$+\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+M\sqrt{L_{2}}\right)\delta G(X_{s},Y_{s})+\frac{2L_{1}}{\varepsilon}\mathbb{E}[G(X_{s},Y_{s})]\rho_{1}(X_{s},Y_{s}).$
Now, from Lemma 3.8 and the fact that $c<\lambda$, we have
@@
$\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+M\sqrt{L_{2}}\right)\delta\int_{0}^{t}e^{cs}\mathbb{E}[G(X_{s},Y_{s})]\mathrm{d}s$
@@
Now, we define $K_{3}=L_{1}K_{1}+2L_{1}/\varepsilon$, $K_{4}=1+2L\varepsilon/\lambda$, and $K_{5}=K_{5}(X_{0},Y_{0})=\varepsilon(\mathbb{E}[V(X_{0})]+\mathbb{E}[V(Y_{0})])$. Then we get
@@
Step 5: In this final step, we establish (3.28), (3.29), and (3.30). We define
@@ 4 Applications
In this section, we explore the applications of the contraction results presented in Section 3. We first show that under the conditions of Theorem 3.4, a propagation of chaos type result holds. Next, we show that under the conditions of Theorem 3.4 or 3.7, (1.1) possesses a unique invariant measure $\mu$ and that starting from any initial distribution $\nu_{0}$, the distribution of $X_{t}$ converges to $\mu$ at an exponential rate.
@@ 4.2 Exponential Ergodicity
As an application of the contraction results, we demonstrate that the McKean-Vlasov process given by (1.1) is exponentially ergodic under the conditions of Theorems 3.4 or 3.7. Let $(X_{t})_{t\geq 0}$ denote the solution to (1.1) with initial condition $X_{0}$. Suppose $\mathcal{L}(X_{0})=\mu$ and denote $P^{*}_{t}\mu=\mathcal{L}(X_{t})$ for $t\geq 0$. We have the semigroup property $P^{*}_{t+s}=P^{*}_{t}P^{*}_{s}$ for any $s,t\geq 0$. As usual, $\mu\in\mathcal{P}(\mathbb{R}^{d})$ is said to be an invariant measure for the semigroup $P_{\cdot}^{*}$ if $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
@@
We adopt the shift coupling idea from the proof of (Wang, , 2018, Theorem 3.1). Let $\delta_{0}$ be the Dirac measure at 00. Then $P_{t}^{*}\delta_{0}=\mathcal{L}(X_{t})$. Next we define for any $s\geq 0$, $(\bar{X}_{t}:=X_{t+s}(0))_{t\geq 0}$. Note that $\bar{X}_{t}$ solves
@@ Appendix A Proofs of Several Lemmas
[…]
@@
First note that Assumption 1.3 gives us $|b(0,\delta_{0})|\leq L_{3}<\infty$, where $\delta_{0}$ is the Dirac measure at 00. Denote $X_{t}=X_{t}(0)$ throughout the proof. Then Itô’s formula gives
@@
$g^{\prime}(s)=\begin{cases}\frac{1}{2}s^{-1/2},&\text{ if }s\geq 1,\\ -\frac{1}{4}s+\frac{3}{4},&\text{ if }0\leq s<1\end{cases}\quad\text{ and }\quad g^{\prime\prime}(s)=\begin{cases}-\frac{1}{4}s^{-3/2},&\text{ if }s\geq 1,\\ -\frac{1}{4},&\text{ if }0\leq s<1.\end{cases}$

### v2

@@ 3.2 Contraction Result Under Dissipative Condition
$-I_{\{r_{t}>0\}}\frac{4\tilde{M}f^{\prime}(r_{t})}{D}\left(|\alpha_{t}^{\top}e_{t}|^{2}\mathrm{rc}^{2}(U_{t})+|\Delta_{t}^{\top}e_{t}|^{2}\mathrm{sc}^{2}(U_{t})\right)\mathrm{d}t$
@@
The last inequality follows because if $\lambda/2\leq 2L\varepsilon$, then $\frac{\lambda}{2}G(X_{t},Y_{t})+\frac{\lambda}{2}-2L\varepsilon\leq-\frac{\lambda}{2}G(X_{t},Y_{t})$, and if $2L\varepsilon\leq\lambda/2$, then
@@
Furthermore, we notice that using the Cauchy-Schwarz inequality, followed by the bound (1.3) and the fact that $\|H\|=1$, we get
@@
$\leq 2\varepsilon f^{\prime}(r_{t})\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|\tilde{M}(|X_{t}|+|Y_{t}|)$
$\leq\frac{\tilde{M}f^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t}),$
@@
$\leq\tilde{M}f^{\prime}(r_{t})\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|G(X_{t},Y_{t})$
$\leq\tilde{M}\sqrt{L_{2}}f^{\prime}(r_{t})\mathrm{sc}^{2}(U_{t})r_{t}G(X_{t},Y_{t})$
$\leq\tilde{M}\sqrt{L_{2}}\delta G(X_{t},Y_{t}).$
@@
$\mathrm{d}[f(r),G(X,Y)]_{t}$ $\leq\frac{\tilde{M}f^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t})\mathrm{d}t+\tilde{M}\sqrt{L_{2}}\delta G(X_{t},Y_{t})\mathrm{d}t.$ (3.44)
@@
$\quad+G(X_{t},Y_{t})\left[\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+\tilde{M}\sqrt{L_{2}}\right)\delta\right]\mathrm{d}t$
$\quad+G(X_{t},Y_{t})\left[-\frac{4\tilde{M}f^{\prime}(r_{t})}{D}\left(|\alpha_{t}^{\top}e_{t}|^{2}\mathrm{rc}^{2}(U_{t})+|\Delta_{t}^{\top}e_{t}|^{2}\mathrm{sc}^{2}(U_{t})\right)\right]\mathrm{d}t$
@@
$\quad+\frac{\tilde{M}f^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t})\mathrm{d}t+\mathrm{d}M_{t}$
@@
$\quad+G(X_{t},Y_{t})\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+\tilde{M}\sqrt{L_{2}}\right)\delta\mathrm{d}t$
@@
$+\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+\tilde{M}\sqrt{L_{2}}\right)\delta G(X_{s},Y_{s})+\frac{2L_{1}}{\varepsilon}\mathbb{E}[G(X_{s},Y_{s})]\rho_{1}(X_{s},Y_{s}).$
We first consider the third term in (3.50). From Lemma 3.8 and the fact that $c<\lambda$, we have
@@
$\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+\tilde{M}\sqrt{L_{2}}\right)\delta\int_{0}^{t}e^{cs}\mathbb{E}[G(X_{s},Y_{s})]\mathrm{d}s$
@@
Next we estimate the second and fourth terms in (3.50). Define $K_{3}=L_{1}K_{1}+2L_{1}/\varepsilon$, $K_{4}=1+2L\varepsilon/\lambda$, and $K_{5}=K_{5}(X_{0},Y_{0})=\varepsilon(\mathbb{E}[V(X_{0})]+\mathbb{E}[V(Y_{0})])$. Then we get
@@
Step 5: In this final step, we establish (3.25), (3.26), and (3.27). First we recall that $c$ is independent of the coupling, and thus does not depend on $\delta$. We define
@@ 4 Applications
In this section, we explore the applications of the contraction results presented in Section 3. We first show that under the conditions of Theorem 3.3, a propagation of chaos type result holds. Next, we show that under the conditions of Theorem 3.3 or 3.6, (1.1) possesses a unique invariant measure $\mu$ and that starting from any initial distribution $\nu_{0}$, the distribution of $X_{t}$ converges to $\mu$ at an exponential rate.
@@ 4.2 Exponential Ergodicity
As an application of the contraction results, we demonstrate that the McKean-Vlasov process given by (1.1) is exponentially ergodic under the conditions of Theorems 3.3 or 3.6. Let $(X_{t})_{t\geq 0}$ denote the solution to (1.1) with initial condition $X_{0}$. Suppose $\mathcal{L}(X_{0})=\mu$ and denote $P^{*}_{t}\mu=\mathcal{L}(X_{t})$ for $t\geq 0$. We have the semigroup property $P^{*}_{t+s}=P^{*}_{t}P^{*}_{s}$ for any $s,t\geq 0$. As usual, $\mu\in\mathcal{P}(\mathbb{R}^{d})$ is said to be an invariant measure for the semigroup $P_{\cdot}^{*}$ if $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
@@
We adopt the shift coupling idea from the proof of (Wang, , 2018, Theorem 3.1). Let $\delta_{0}$ be the Dirac measure at 0. Then $P_{t}^{*}\delta_{0}=\mathcal{L}(X_{t})$. Next we define for any $s\geq 0$, $(\bar{X}_{t}:=X_{t+s}(0))_{t\geq 0}$. Note that $\bar{X}_{t}$ solves
@@ Appendix A Proofs of Several Lemmas
$\int_{0}^{t}I_{\{|Z_{s}|^{2}<\varepsilon\}}\bigg(\bigg(-\frac{r_{s}^{2}}{4\varepsilon^{3/2}}+\frac{3}{4\varepsilon^{1/2}}\bigg)\left(\langle Z_{s},\beta_{s}\rangle+\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})\right)-\frac{1}{4\varepsilon^{3/2}}|\Delta_{s}^{\top}Z_{s}|^{2}\bigg)\mathrm{d}s\to 0,\ \ \ \text{a.s.}$ (A.8)
@@
First note that Assumption 1.3 gives us $|b(0,\delta_{0})|\leq L_{3}<\infty$, where $\delta_{0}$ is the Dirac measure at 0. Denote $X_{t}=X_{t}(0)$ throughout the proof. Then Itô’s formula gives
@@
[…]